News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 30 May 2003 7.00am EDT

Last week the observing schedule was halted due to high radiation associated with solar flare activity. The loads were halted on May 29 at 4:47pm EDT through a ground command to execute the SI Safing SCS 107. This ensured that the accumulated radiation dose for ACIS remained below the allowed threshold. Loads will be re-started once radiation levels reach an acceptable level.

A real-time procedure was run on 27 May to lower the ACA CCD temperature from -11 to -12 deg C to mitigate the increasing number of warm pixels. This is the second of five planned 1 degree step reductions. An assessment of aspect camera data following the first reduction on May 8 shows an expected improvement.

The IRU-1 gyro-1 current remained quiet and near 140 mA during the last week (a slight decrease in the current level).

Two Chandra press releases were issued on May 26 in conjunction with the AAS meeting. The first discussing the oxygen-rich Supernova remnant 0103-72.6, and the second describing a new way to determine if a young star is surrounded by a planet-forming disk like our early Sun. For details see:

The observing schedule for next week is shown below and includes an observation of 3C279 coordinated with Integral.

Radiation Belts                      Jun 1
SGR 1900+14            ACIS-S        Jun 2
FIRST J164442+261913   ACIS-S
3C 279                 HRC-S/LETG
F15130-1958            ACIS-S
HR 5512                HRC-I
W51B                   ACIS-I
4C 74.26               ACIS-S/HETG   Jun 3
Radiation Belts                      Jun 4
4U 1626-67             ACIS-S/HETG
RCS0439.6-2905         ACIS-S        Jun 6
XMM 1HR - Field 3      ACIS-I
Radiation Belts
HIP 92680              ACIS-S/HETG   Jun 7
XMM 1Hr - Field 4      ACIS-I        Jun 8


All spacecraft subsystems, with the exception of the MUPS thruster A-1 under investigation, continued to operate nominally.

Return to Status Reports